OpenStudy (anonymous):

Solve 5^(2x) = 7^(x−1). A. 1.52864 B −1 C −1.52864 D 1

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 5^{2x} =7^{x-1}}\) \(\color{#000000 }{ \displaystyle \ln(5^{2x} )=\ln(7^{x-1})}\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 2x\ln(5)=(x-1)\ln(7)}\)

OpenStudy (solomonzelman):

etc....

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

If that is it, then what was the problem solving it from the beginning?